1. (5.) Find original function and reduced form f(x) = 3(x-1) (x+2) / (x-4) (x+2)
1
Expert's answer
2021-09-13T00:26:26-0400

Original function


"f(x)=\dfrac{3(x-1)(x+2)}{(x-4)(x+2)}, x\not=-2, x\not=4"

Reduced form


"f(x)=\dfrac{3(x-1)}{x-4}, x\not=-2, x\not=4"